Use the distance formulas to decide whether line JKL approximately equals line KL.

J (3,-5)

K (-1,2)

L (-5,-5)

thank you

Answer by Alan3354(69443)

Use the distance formulas to decide whether line JKL approximately equals line KL.
J (3,-5)
K (-1,2)
L (-5,-5)
---------
JK: s^2 = diffy^2 + diffx^2
s^2 = (-5-2)^2 + (-1-3)^2
s^2 = 49 + 16 = 65
------------
KL: s^2 = 16 + 49 = 65
Same length, not approximate